# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ III Approximate PTC rigidity for local distributions
Result 2.(Relaxation of the local set $\mathcal{L}_{\triangle}^{2}$) Any triangle local distribution $P(a,b,c)\in\mathcal{L}_{\triangle}^{2}$ that is close to the PTC slice of the probability set $\delta(P,\mathcal{P}_{PTC}^{2})\leq\varepsilon$ is also close to the local PTC set $\delta(P,\mathcal{L}^{2}_{\triangle|PTC})\leq\text{f}_{\varepsilon}$ (see Eq. 16).
Figure 2: A graphical representation of the construction used in the proof of approximate PTC rigidity. The coordinates along the axes represent different values of the local variables $\beta$ and $\gamma$, here discredited to 15 possible values each. The color of each square gives the value of Alice’s outcome $a(\beta,\gamma)=0$ (yellow) and $a(\beta,\gamma)=1$ (blue). On the left, the lighter region depicts the set $\Upsilon$ on which the binary token functions $t_{\beta}$ and $t_{\gamma}$ are defined (given below the axes) and satisfy $a(\beta,\gamma)=t_{\beta}\oplus t_{\gamma}\oplus 1$. On the right, the token functions are prolonged on the whole set of local variables by setting them to 0 outside of $\Upsilon$. The values obtained by applying the PTC response function $a(\beta,\gamma)=t_{\beta}\oplus t_{\gamma}\oplus 1$ (right) are not guaranteed to match the observed values $a(\beta,\gamma)$ (left) outside of $\Upsilon$.
In words, the approximate rigidity allows us to define a cone of distributions ^$\widehat{\mathcal{L}}^{2}_{\triangle}$ that are accessible when going away from the PTC slice. Loosely speaking – if we only move a distance $\varepsilon$ from $\mathcal{L}_{\triangle|PTC}^{2}$ perpendicular to the PTC slice we can not move more than a distance $\varepsilon$ inside the slice. Any distribution outside of the cone is guaranteed to be nonlocal, i.e. we have established an outer approximation of the local set $\mathcal{L}_{\triangle}^{2}\subset\widehat{\mathcal{L}}^{2}_{\triangle}$. It is an open question whether all quantum distributions are also contained in this cone.
@@
IV Noise-robust nonlocality in $\mathcal{Q}^{4}_{\triangle}$
@@
|$\ket{\psi}=\lambda_{0}\ket{01}+\lambda_{1}\ket{10},$ (23)
with real and positive $\lambda_{0}^{2}+\lambda_{1}^{2}=1$, to the neighbouring parties. While each party performs the same projective measurement $\{\Pi_{x_{1},x_{2}}=\outerproduct{\phi_{x_{1},x_{2}}}{\phi_{x_{1},x_{2}}}\}$ | } on the two qubits it receives from neighboring sources, with the POVM elements given by projectors on the states
@@
with real positive $u,v,w,z$ satisfying $u^{2}+v^{2}=w^{2}+z^{2}=1$. The resulting family of quantum distribution is termed “tilted Bell state measurements” (TBSM) distribution, since the local measurements can be viewed as tilted Bell state measurements.
$P(1,i,1,j,1,k)$ $=(\lambda_{0}^{3}u_{i}u_{j}u_{k}+\lambda_{1}^{3}v_{i}v_{j}v_{k})^{2}$ (25)
$P(1,i,0,j,0,k)$ $=(\lambda_{0}\lambda_{1}^{2}u_{i}w_{j}z_{k}+\lambda_{1}\lambda_{0}^{2}v_{i}w_{k}z_{j})^{2}\,\circlearrowleft$ (26)
where $u_{0}=u,\,v_{0}=v\,,w_{0}=w\,,z_{0}=z,\,u_{1}=v,\,v_{1}=-u,\,w_{1}=z$ and $z_{1}=-w$
By setting $w=1-z=0$ one recovers the RGB4 family of distributions [14], while merging the outcomes $(0,0)$ and $(0,1)$ leads to the three-output family of quantum distributions introduced in [22].
As a warm-up, we now present a proof of nonlocality of the distributions in Eq. (25) for a subset of parameter values $(\lambda_{0},u,m)$. It has the advantage of streamlining the proof techniques used in [14] and subsequently [22], while relying on the same basic idea, This will be helpful when we move to the noisy case.
IV.1 Nonlocality of the noiseless TBSM distributions
@@
satisfy the parity condition $a_{1}\oplus b_{1}\oplus c_{1}=1$. By rigidity, it follows that for any underlying local model there exist the binary token functions $T_{\xi}$ that explain the distribution
@@
where we introduced the compact notation $\bm{x}_{i}=(a_{i},b_{i},c_{i})$, $\bm{t}=(t_{\alpha},t_{\beta},t_{\gamma})$ and
$P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})=P_{\scriptscriptstyle ptc}(a_{1}|t_{\beta},t_{\gamma})P_{\scriptscriptstyle ptc}(b_{1}|t_{\alpha},t_{\gamma})P_{\scriptscriptstyle ptc}(c_{1}|t_{\alpha},t_{\beta}),$ (29)
with the response functions of Eq. (19). Furthermore, if $E_{a_{1}},E_{b_{1}},E_{c_{1}}\neq 0$ the token distribution $p(\bm{t})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is known exactly via Eq. (5).
This conclusion is of course independent of whether we discard the outcomes $\bm{x}_{2}$ or not. Hence, for any local model resulting in the TBSM distribution, we can define the random variable $\bm{t}$, indicating the behavior of the tokens, using the Chain rule we have $P(\bm{x}_{1},\bm{x}_{2})=\sum_{\bm{t}}p(\bm{t})P(\bm{x}_{2}|\bm{t})P(\bm{x}_{1}|\bm{t},\bm{x}_{2})$. From the rigidity, we can infer that $\bm{x}_{1}$ is solely a function of $\bm{t}$, leading to
@@
Here, all the distributions $P(\bm{x}_{2}|\bm{t})\in\mathcal{L}_{\triangle}^{2}$ must be triangle local. Furthermore, these distributions are not independent, since the output of a party (e.g. $a_{2}$) can not be influenced by the source that is not connected to it (that is by $\alpha$ and the token value $t_{\alpha}$). This observation implies the following constraints on the marginal distributions $P(a_{2}|\bm{t})=\sum_{b_{2},c_{2}}P(\bm{x}_{2}|\bm{t})$
@@
Result 3. (Nonlocality of TBSM distributions) For any triangle local distribution $P(\bm{x}_{1},\bm{x}_{2})\in\mathcal{L}_{\triangle}^{4}$ satisfying $a_{1}\oplus b_{1}\oplus c_{1}=1$ , there exist eight probability distributions $P(\bm{x}_{2}|\bm{t})$ such that
@@
where $P_{\scriptscriptstyle PTC}$ is the parity token counting response function and the token distribution $p(\bm{t})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is uniquely determined from $P(\bm{x}_{1})$ via Eq. (5) (except degenerate cases were it is not unique).
If we were able to show that these conditions can not be satisfied together, we could conclude that the distribution $P(\bm{x}_{1},\bm{x}_{2})$ is nonlocal. Yet, the locality condition $(3.iii)$ here is very challenging to formalize since we do not have a good characterization of the local set $\mathcal{L}_{\triangle}^{2}$. A simple solution is to relax it and keep the other two conditions. This relaxation gives rise to a simple linear program (LP) that must be feasible for any triangle local $P(\bm{x}_{1},\bm{x}_{2})$ with $a_{1}\oplus b_{1}\oplus c_{1}=1$, that we give in appendix C.
@@
In Fig. 3, we present the results of the LP for $\lambda_{0}^{2}=0.22$. One notes that Result 3 can be used to prove the nonlocality of the noisy distribution, if the noise does not perturb the parity of the $\bm{x}_{1}$ outcomes. For example, the local dephasing noise given by the channel $\mathcal{E}_{d}:\outerproduct{\psi}{\psi}\mapsto\varrho$ with
@@
is of this type. In the appendix C we give the LP which finds the minimal amount of dephasing noise such that the distribution becomes triangle local. Its results are presented in Fig. 4 for $\lambda_{0}^{2}=0.22$.
@@
Nevertheless, this still only applies to a slice of the quantum set $\mathcal{Q}_{\triangle}^{4}$. In the next section, we derive an approach that singles out the nonlocality of a full-measure subset of triangle quantum distributions. Interestingly, we observe that the largest robustness to white noise, as well as for total-variation distance, is obtained for the same parameters $\lambda_{0}^{2}=0.22,\phi_{u}=0.42$.
IV.2 Nonlocality of the noisy TBSM distributions
Our main idea now is pretty simple – adjust the LP-based nonlocality proof of the previous section to the setting of approximate PTC rigidity discussed in Section III.
Indeed, let us consider any triangle local distribution
$P(\bm{x}_{1},\bm{x}_{2})=\mathds{E}\big{(}P_{A}(a_{1},a_{2}|\beta,\gamma)P_{B}(b_{1},b_{2}|\alpha,\gamma)P_{C}(\dots)\big{)}$ (34)
@@ IV.2 Nonlocality of the noisy TBSM distributions
$\widetilde{P}(\bm{x}_{1},\bm{x}_{2})=\mathds{E}\big{(}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P_{A}(a_{2}|\beta,\gamma)P_{B}(b_{2}|\alpha,\gamma)\dots).$ (35)
The distribution $\widetilde{P}$ is triangle local, by construction it agrees with the initial distribution on outcomes $\bm{x}_{2}$ ($\widetilde{P}(\bm{x}_{2})=P(\bm{x}_{2})$), and is PTC local for the outcomes $\bm{x}_{1}$. Furthermore, we know that the two distributions match $P(\bm{x}_{1},\bm{x}_{2},\bm{\xi}\in\Upsilon)=\widetilde{P}(\bm{x}_{1},\bm{x}_{2},\bm{\xi}\in\Upsilon)$ on the subset $\Upsilon$ of local variable values with $\text{Pr}(\bm{\xi}\in\Upsilon)\geq 1-\text{f}_{\varepsilon}$, which is equivalent to a bound on their total variation distance
@@
[…]
@@
In turn, the distribution $\widetilde{P}$ satisfies the premises of the Result 3. We could thus verify the nonlocality of $P$ by showing that no distribution $\widetilde{P}$ compatible with both Result 3 and 4 exists. A complication here is that the relation (5) between the distribution $\widetilde{P}(\bm{x}_{1})$ and the underlying token distribution $p(\bm{t})$ is non-linear. It is thus easier to introduce the token distributions $p_{\alpha},p_{\beta},p_{\gamma}$ as additional variables rather than computing them from $\widetilde{P}$. That way we can combine results 3 and 4 into the following corollary.
[…]
(*$(*)\,\widetilde{P}(\bm{x}_{1},\bm{x}_{2})=\sum_{\bm{t}}p(\bm{t})P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2}|\bm{t})$ (38)
@@
Figure 5: White Noise Robustness. Maximal amount of white noise below which we can prove the nonlocality of TBSM distributions. The largest amount of noise that can be tolerated is here $\omega=0.5\%$; note that this value can be improved slightly to $\omega=0.544\%$ (see Appendix. D.1). Parameters: $u=\cos(\phi_{u})$, $w=\cos(\phi_{w})$ and $\lambda_{0}^{2}=0.22$.
As in the ideal case to verify this feasibility problem we start by forgetting the condition $(3.iii)$ requiring the distributions $P(\bm{x}_{2}|\bm{t})$ to be local. Note also that the condition $(4.ii)$ on the total variation distance $\delta(P,\widetilde{P})\leq\text{f}_{\varepsilon}$ is equivalent to the existence of two probability distributions $P^{\prime}(\bm{x}_{1},\bm{x}_{2})$ and $\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})$ fulfilling $P+\text{f}_{\varepsilon}P^{\prime}=\widetilde{P}+\text{f}_{\varepsilon}\widetilde{P}$. By introducing these probabilities as variables we can rewire the conditions $(4.iii)$ and (*)(*)( * ) as
@@
$\begin{split}\frac{1}{2}&\left(1+\sqrt{\frac{(|E_{b_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon))(|E_{c_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon))}{|E_{a_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon)}}\right)\leq p_{\alpha}(0)\\ &\leq\frac{1}{2}\left(1+\sqrt{\frac{(|E_{b_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon))(|E_{c_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon))}{|E_{a_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon)}}\right)\quad\circlearrowleft.\end{split}$ (42)
In the Appendix D we show how using this bound the feasibility problem in Eq. (38) can be relaxed to an LP, or a tighter collection of LPs for a grid of values $p_{\xi}(0)$. To illustrate this approach we now consider the family of noisy TBSM distributions, where local white noise is added to the pure state in Eq. (23) leading to the density operator
$\rho=(1-\omega)\outerproduct{\psi}{\psi}+\omega\frac{\mathds{1}}{2}\otimes\frac{\mathds{1}}{2}.$ (43)
The resulting distribution satisfies $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)=\frac{1}{2}\left(1+(1-\omega)^{3}\right)$. In Fig. 5 we plot the maximal amount of white noise $\omega$ for which we can prove the nonlocality of the distribution for $\lambda_{0}^{2}=0.22$ and different measurement parameters. The maximal value that we find is about $\omega=0.67\%$ at $\phi_{u}=0.42$.
@@
Figure 6: Total Variation. The size of the total variation distance ball $\mathcal{B}_{\varepsilon}(\bar{P})$ in Eq. (44) around the ideal TBSM distributions, for which we prove nonlocality. The maximal value we find is $\varepsilon=0.22\%$, which can be improved to $\varepsilon=0.24\%$ (see Appendix. D.1). Parameters: $u=\cos(\phi_{u})$, $w=\cos(\phi_{w})$ and $\lambda_{0}^{2}=0.22$.
IV.3 Nonlocality of regions of the quantum set.
So far we argued how the nonlocality of a given distribution $P(\bm{x}_{1},\bm{x}_{2})$ can be demonstrated. However, to better understand the separation between the quantum and the local sets it is more insightful to derive expressions that single out whole regions of the quantum sets that are guaranteed to be nonlocal, similarly to the usual Bell tests. In the case of nonlocality without inputs, considered here, such expression must be nonlinear. We now derive such ”Bell tests” in terms of total variation distance balls around a target quantum distribution.
@@
In addition, if the ball contains some triangle local distribution $P$ by Result 4 there must exist a local PTC distribution $\widetilde{P}$ with $\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),P(\bm{x}_{1},\bm{x}_{2}))\leq\text{f}_{\varepsilon}$ by (4.ii) and $\delta(\widetilde{P}(\bm{x}_{2}),P(\bm{x}_{2}))=0$ by (4.iii). By the triangle inequality, $\widetilde{P}$ must also satisfy
$\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),\bar{P}(\bm{x}_{1},\bm{x}_{2}))$ $\leq\delta(\widetilde{P},P)+\delta(P,\bar{P})\leq\text{f}_{\varepsilon}+\varepsilon$
@@
Result 5. (Nonlocality of a ball $B_{\varepsilon}(\bar{P})$) If the ball $B_{\varepsilon}(\bar{P})\subset\mathcal{P}_{\triangle}^{4}$ contains triangle local distributions, there exists a triangle local distribution $\widetilde{P}(\bm{x}_{1},\bm{x}_{2})$ such that

### v3

@@ 4 Approximate PTC rigidity for local distributions
Result 2. (Relaxation of the local set $\mathcal{L}_{\triangle}^{2}$) Any triangle-local distribution $P(a,b,c)\in\mathcal{L}_{\triangle}^{2}$ that is close to the PTC slice of the probability set $\delta(P,\mathcal{P}_{PTC}^{2})\leq\varepsilon$ is also close to the local (also triangle-local) PTC set $\delta(P,\mathcal{L}^{2}_{\triangle|PTC})\leq\text{f}_{\varepsilon}$ (see Eq. 15).
Figure 3: A graphical representation of the construction used in the proof of approximate PTC rigidity. The coordinates along the axes represent different values of the local variables $\beta$ and $\gamma$, here discretized to 15 possible values each. The color of each square gives the value of Alice’s outcome $a(\beta,\gamma)=0$ (yellow) and $a(\beta,\gamma)=1$ (blue). On the left, the lighter region depicts the set $\Upsilon$ on which the binary token functions $t_{\beta}$ and $t_{\gamma}$ are defined (given below the axes) and satisfy $a(\beta,\gamma)=t_{\beta}\oplus t_{\gamma}\oplus 1$. On the right, the token functions are extended on the whole set of local variables by setting them to 0 outside of $\Upsilon$. The values obtained by applying the PTC response function $a(\beta,\gamma)=t_{\beta}\oplus t_{\gamma}\oplus 1$ (right) are not guaranteed to match the observed values $a(\beta,\gamma)$ (left) outside of $\Upsilon$.
In words, the approximate rigidity allows us to define a cone of distributions $\widehat{\mathcal{L}}^{2}_{\triangle}$ that are accessible when going away from the PTC slice, as illustrated in Fig. 4. Loosely speaking – it contains triangle-local distributions at a distance $\varepsilon$ to the PTC slice that must also be close enough to the local PTC set such that we can not move more than a distance $\varepsilon$ inside the slice. Any distribution outside of the cone is guaranteed to be nonlocal, i.e. we have established an outer approximation of the local set $\mathcal{L}_{\triangle}^{2}\subset\widehat{\mathcal{L}}^{2}_{\triangle}$. It is an open question whether all quantum distributions are also contained in this cone.
@@
Figure 4: A graphical representation of Result 2. Consider any triangle-local distribution $P(a,b,c)\in\mathcal{L}^{2}_{\triangle}$ which is at a distance $\varepsilon=\delta(P,\mathcal{P}_{PTC}^{2})$ to the PTC slice containing all probability distributions satisfying $\text{Pr}(a\oplus b\oplus c\neq 1)=0$. Result 2 implies that $P(a,b,c)$ must also be ”close enough” $\delta(P,\mathcal{L}_{\triangle|PTC}^{2})\leq\text{f}_{\varepsilon}$ to the local PTC set $\mathcal{L}^{2}_{\triangle|PTC}$ of which we have full characterization. Here, $\delta$ is the total variation distance, and the function $\text{f}_{\varepsilon}\leq 3\varepsilon$ is defined in Eq. (15).
5 Noise-robust nonlocality in $\mathcal{Q}^{4}_{\triangle}$
@@
[…]
$\begin{split}P(1,i,1,j,1,k)&=(\lambda_{0}^{3}u_{i}u_{j}u_{k}+\lambda_{1}^{3}v_{i}v_{j}v_{k})^{2}\\ P(1,i,0,j,0,k)&=(\lambda_{0}\lambda_{1}^{2}u_{i}w_{j}z_{k}+\lambda_{1}\lambda_{0}^{2}v_{i}w_{k}z_{j})^{2}\,\circlearrowleft\end{split}$ (24)
[…]
By setting w=1z=0w=1-z=0 one recovers the RGB4 family of distributions [16], while merging the outcomes $(0,0)$ and $(0,1)$ leads to the three-output family of quantum distributions introduced in [20].
As a warm-up, we now present a proof of nonlocality of the distributions in Eq. (24) for a subset of parameter values $(\lambda_{0},u,w)$. It has the advantage of streamlining the proof techniques used in [16] and subsequently [20], while relying on the same basic idea. This will be helpful when we move to the noisy case.
5.1 Nonlocality of the noiseless TBSM distributions
@@
satisfy the parity condition $a_{1}\oplus b_{1}\oplus c_{1}=1$. By rigidity, it follows that for any underlying local model there exist the binary token values $t_{\alpha},t_{\beta},t_{\gamma}$ that explain the distribution
@@
where $\mathds{E}$ is the expected value over $\bm{t}$ and we introduced the compact notation $\bm{x}_{i}=(a_{i},b_{i},c_{i})$, $\bm{t}=(t_{\alpha},t_{\beta},t_{\gamma})$ and $P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})=P_{\scriptscriptstyle ptc}(a_{1}|t_{\beta},t_{\gamma})P_{\scriptscriptstyle ptc}(b_{1}|t_{\alpha},t_{\gamma})P_{\scriptscriptstyle ptc}(c_{1}|t_{\alpha},t_{\beta}),$ with the response functions of Eq. (18). Furthermore, if $E_{a_{1}},E_{b_{1}},E_{c_{1}}\neq 0$ the token distribution $p(\bm{t})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is known exactly via Eq. (5).
This conclusion is, of course, independent of whether we discard the outcomes $\bm{x}_{2}$ or not. Hence, for any local model resulting in the TBSM distribution, we can define the random variable $\bm{t}$, indicating the behavior of the tokens. Using the chain rule, we have $P(\bm{x}_{1},\bm{x}_{2})=\sum_{\bm{t}}p(\bm{t})P(\bm{x}_{2}|\bm{t})P(\bm{x}_{1}|\bm{t},\bm{x}_{2})$. From the rigidity, we can infer that $\bm{x}_{1}$ is solely a function of $\bm{t}$, leading to
@@
Here, all the distributions $P(\bm{x}_{2}|\bm{t})\in\mathcal{L}_{\triangle}^{2}$ must be triangle-local. Furthermore, these distributions are not independent, since the output of a party (e.g. $a_{2}$) can not be influenced by the source that is not connected to it (that is by $\alpha$ and the token value $t_{\alpha}$). This observation implies the following constraints on the marginal distributions $P(a_{2}|\bm{t})=\sum_{b_{2},c_{2}}P(\bm{x}_{2}|\bm{t})$
@@
Result 3. (Nonlocality of TBSM distributions) For any triangle-local distribution $P(\bm{x}_{1},\bm{x}_{2})\in\mathcal{L}_{\triangle}^{4}$ satisfying $a_{1}\oplus b_{1}\oplus c_{1}=1$ , there exist eight probability distributions $P(\bm{x}_{2}|\bm{t})$ such that
@@
where $P_{\scriptscriptstyle PTC}$ is the parity token counting response function and the token distribution $p(\bm{t})=p_{\alpha}(t_{\alpha})p_{\beta}(t_{\beta})p_{\gamma}(t_{\gamma})$ is uniquely determined from $P(\bm{x}_{1})$ via Eq. (5) (except degenerate cases where it is not unique).
If we were able to show that these conditions cannot be satisfied together, we could conclude that the distribution $P(\bm{x}_{1},\bm{x}_{2})$ is nonlocal. Yet, the locality condition $(3.iii)$ here is very challenging to formalize since we do not have a good characterization of the local set $\mathcal{L}_{\triangle}^{2}$. A simple solution is to relax it and keep the other two conditions. This relaxation gives rise to a simple linear program (LP) that must be feasible for any triangle-local $P(\bm{x}_{1},\bm{x}_{2})$ with $a_{1}\oplus b_{1}\oplus c_{1}=1$, that we give in Appendix C.
@@
[…]
is of this type. In Appendix C we give the LP which finds the minimal amount of dephasing noise such that the distribution becomes triangle-local. Its results are presented in Fig. 6 for $\lambda_{0}^{2}=0.22$.
@@
Nevertheless, this still only applies to a slice of the quantum set $\mathcal{Q}_{\triangle}^{4}$. In the next Section, we derive an approach that singles out the nonlocality of a full-measure subset of triangle quantum distributions. Interestingly, we observe that the largest robustness to white noise, as well as for total-variation distance, is obtained for the same parameters $\lambda_{0}^{2}=0.22,\phi_{u}=0.42,\phi_{w}=0$, with $u=\cos(\phi_{u})$ and $w=\cos(\phi_{w})$.
5.2 Nonlocality of the noisy TBSM distributions
Our main idea now is pretty simple – adjust the LP-based nonlocality proof of the previous Section to the setting of approximate PTC rigidity discussed in Section 4.
Indeed, let us consider any triangle-local distribution
$P(\bm{x}_{1},\bm{x}_{2})$
$=\mathds{E}\big{(}P_{A}(a_{1},a_{2}|\beta,\gamma)P_{B}(b_{1},b_{2}|\alpha,\gamma)P_{C}(\dots)\big{)}$
@@ 5.2 Nonlocality of the noisy TBSM distributions
$\widetilde{P}(\bm{x}_{1},\bm{x}_{2})$ (31)
$=\mathds{E}\big{(}P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P_{A}(a_{2}|\beta,\gamma)P_{B}(b_{2}|\alpha,\gamma)\dots).$
The distribution $\widetilde{P}$ is triangle-local, by construction it agrees with the initial distribution on outcomes $\bm{x}_{2}$ ($\widetilde{P}(\bm{x}_{2})=P(\bm{x}_{2})$), and is PTC local for the outcomes $\bm{x}_{1}$. Furthermore, we know that the two distributions match $P(\bm{x}_{1},\bm{x}_{2};\bm{\xi}\in\Upsilon)=\widetilde{P}(\bm{x}_{1},\bm{x}_{2};\bm{\xi}\in\Upsilon)$ on the subset $\Upsilon$ of local variable values with $\text{Pr}(\bm{\xi}\in\Upsilon)\geq 1-\text{f}_{\varepsilon}$, which is equivalent to a bound on their total variation distance
@@
Result 4. (Nonlocality of noisy TBSM distribution) For any triangle-local distribution $P(\bm{x}_{1},\bm{x}_{2})\in\mathcal{L}_{\triangle}^{4}$ satisfying $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)\geq 1-\varepsilon$, there exists a triangle-local distribution $\widetilde{P}(\bm{x}_{1},\bm{x}_{2})$ such that
@@
In turn, the distribution $\widetilde{P}$ satisfies the premises of Result 3. We could thus verify the nonlocality of $P$ by showing that no distribution $\widetilde{P}$ compatible with both Result 3 and 4 exists. A complication here is that the relation (5) between the distribution $\widetilde{P}(\bm{x}_{1})$ and the underlying token distribution $p(\bm{t})$ is nonlinear. It is thus easier to introduce the token distributions $p_{\alpha},p_{\beta},p_{\gamma}$ as additional variables rather than computing them from $\widetilde{P}$. That way we can combine results 3 and 4 into the following corollary.
Corollary of Result 4. For any triangle-local distribution $P(\bm{x}_{1},\bm{x}_{2})\in\mathcal{L}_{\triangle}^{4}$ satisfying $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)\geq 1-\varepsilon$, there exist probability distributions $P(\bm{x}_{2}|\bm{t})$, $p_{\xi}(t_{\xi})$ and $\widetilde{P}(\bm{x}_{1},\bm{x}_{2})$ such that $(4.ii),(4.iii),(3.ii),(3.iii)$ and
$(*)\,\widetilde{P}(\bm{x}_{1},\bm{x}_{2})=\sum_{\bm{t}}p(\bm{t})P_{\scriptscriptstyle PTC}(\bm{x}_{1}|\bm{t})P(\bm{x}_{2}|\bm{t})$ (34)
@@
Figure 7: White Noise Robustness. Maximal amount of white noise below which we can prove the nonlocality of TBSM distributions. The largest amount of noise that can be tolerated is here $\omega=0.5\%$; note that this value can be improved slightly to $\omega=0.544\%$ (see Appendix D.1). Parameters: $u=\cos(\phi_{u})$, $w=\cos(\phi_{w})$ and $\lambda_{0}^{2}=0.22$.
As in the ideal case to verify this feasibility problem we start by forgetting the condition $(3.iii)$ requiring the distributions $P(\bm{x}_{2}|\bm{t})$ to be local. Note also that the condition $(4.ii)$ on the total variation distance $\delta(P,\widetilde{P})\leq\text{f}_{\varepsilon}$ is equivalent to the existence of two probability distributions $P^{\prime}(\bm{x}_{1},\bm{x}_{2})$ and $\widetilde{P}^{\prime}(\bm{x}_{1},\bm{x}_{2})$ fulfilling $P+\text{f}_{\varepsilon}P^{\prime}=\widetilde{P}+\text{f}_{\varepsilon}\widetilde{P}^{\prime}$ (see Lemma 1 in Appendix D). By introducing these probabilities as variables we can rewire the conditions $(4.iii)$ and ()(*) as
@@
$\begin{split}\frac{1}{2}&\left(1+\sqrt{\frac{(|E_{b_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon))(|E_{c_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon))}{|E_{a_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon)}}\right)\\ &\leq p_{\alpha}(0)\leq\\ \frac{1}{2}&\left(1+\sqrt{\frac{(|E_{b_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon))(|E_{c_{1}}^{\Lambda}|+(2\text{f}_{\varepsilon}-\varepsilon))}{|E_{a_{1}}^{\Lambda}|-(2\text{f}_{\varepsilon}-\varepsilon)}}\right),\end{split}$ (38)
up to cyclic permutations.
In Appendix D we show how using this bound the feasibility problem in Eq. (34) can be relaxed to an LP, or a tighter collection of LPs for a grid of values $p_{\xi}(0)$. To illustrate this approach we now consider the family of noisy TBSM distributions, where local white noise is added to the pure state in Eq. (22) leading to the density operator
[…]
The resulting distribution satisfies $\text{Pr}(a_{1}\oplus b_{1}\oplus c_{1}=1)=\frac{1}{2}\left(1+(1-\omega)^{3}\right)$. In Fig. 7 we plot the maximal amount of white noise $\omega$ for which we can prove the nonlocality of the distribution for $\lambda_{0}^{2}=0.22$ and different measurement parameters. The maximal value that we find is about $\omega=0.55\%$ at $\phi_{u}=0.42$.
@@
Figure 8: Total Variation. The size of the total variation distance ball $\mathcal{B}_{\varepsilon}(\bar{P})$ in Eq. (40) around the ideal TBSM distributions, for which we prove nonlocality. The maximal value in the plot is $\varepsilon=0.22\%$, which can be improved to $\varepsilon=0.24\%$ (see Appendix D.1). Parameters: $u=\cos(\phi_{u})$, $w=\cos(\phi_{w})$ and $\lambda_{0}^{2}=0.22$.
5.3 Nonlocality of regions of the quantum set.
So far we argued how the nonlocality of a given distribution $P(\bm{x}_{1},\bm{x}_{2})$ can be demonstrated. However, to better understand the separation between the quantum and the local sets it is more insightful to derive expressions that single out whole regions of the quantum sets that are guaranteed to be nonlocal, similarly to the usual Bell tests. In the case of nonlocality without inputs, considered here, such an expression must be nonlinear. We now derive such ”Bell tests” in terms of total variation distance balls around a target quantum distribution.
@@
In addition, if the ball contains some triangle-local distribution $P$ by Result 4 there must exist a local PTC distribution $\widetilde{P}$ with $\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),P(\bm{x}_{1},\bm{x}_{2}))\leq\text{f}_{\varepsilon}$ by (4.ii) and $\delta(\widetilde{P}(\bm{x}_{2}),P(\bm{x}_{2}))=0$ by (4.iii). By the triangle inequality, $\widetilde{P}$ must also satisfy
$\delta(\widetilde{P}(\bm{x}_{1},\bm{x}_{2}),\bar{P}(\bm{x}_{1},\bm{x}_{2}))$ $\leq\delta(\widetilde{P},P)+\delta(P,\bar{P})$
$\leq\text{f}_{\varepsilon}+\varepsilon$
@@
Result 5. (Nonlocality of a ball $B_{\varepsilon}(\bar{P})$) If the ball $B_{\varepsilon}(\bar{P})\subset\mathcal{P}_{\triangle}^{4}$ contains triangle-local distributions, there exists a triangle-local distribution $\widetilde{P}(\bm{x}_{1},\bm{x}_{2})$ such that